OpenStudy (anonymous):

Put parabola in Standard form: X^2+4x-4y+16=0

OpenStudy (anonymous):

y = 1/4x^2+x+4 If I'm not mistaken

OpenStudy (campbell_st):

or (x + 2)^2 = 4 ( y - 3)